# Model-Based Reward Shaping for Adversarial Inverse Reinforcement Learning in Stochastic Environments

### v1

@@ Appendix C Performance Gap Analysis
Lemma C.2 (Explicit Feasible Reward Function (Metelli et al., 2021)).
@@
With  Eq. 16, we can derive the following error bound between $R\in\mathcal{R}^{E}_{\mathfrak{B}}$ and $\hat{R}^{E}\in\mathcal{R}_{\hat{\mathfrak{B}}}$.
Theorem C.3 (Reward Function Error Bound).
@@
Proof.
From  Lem. C.2, we can derive the following representations of $R$ and $\hat{R}$ with the same set of $V$ and $\xi$:
@@
Next, we will propagate this bound to the value functions of optimal policy regarding different reward functions $R$ and $\hat{R}$. From the traditional Value iteration, we can write out the value function.
$V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\mathcal{T}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right].$ (18)
Lemma C.4 (Value Function Error under same policy and different rewards and MDP).
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||$: the performance difference of the same policy in different MDPs.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|.$ (19)
Proof.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||$ (20)
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}||\mathcal{T}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]$
$+\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[\hat{R}(s,a)+\gamma V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})\right]||$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}(\epsilon_{\mathcal{T}}\frac{R_{\max}}{1-\gamma}+\hat{\mathcal{T}}(s^{\prime}|s,a)\epsilon_{\mathcal{T}}(\frac{\gamma R_{\max}}{1-\gamma}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||))$
$=\sum_{a\in\mathcal{A}}\pi(a|s)(\epsilon_{\mathcal{T}}\frac{R_{\max}}{1-\gamma}|\mathcal{S}|+\epsilon_{\mathcal{T}}\frac{\gamma R_{\max}}{1-\gamma}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||)$
$\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{1-\gamma}R_{\max}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||$
@@
Lemma C.5.
Let $\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{2}}(s)\|$ denote the performance difference between different policies $\pi_{1}$ and $\pi_{2}$ in the same learned MDP (Viano et al., 2021; Zhang et al., 2020). The following inequality holds:
$\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{2}}(s)\|\leq\frac{\gamma}{(1-\gamma)^{2}}\epsilon_{\mathcal{T}}R_{\max}.$
Theorem C.6 (Performance Difference Bound).
The performance difference between the optimal policies ($\pi^{*}$ and $\hat{\pi}^{*}$) in corresponding MDPs ($\mathcal{M}^{\prime}\cup R$ and ^$\hat{\mathcal{M}}^{\prime}\cup\hat{R}$) can be bounded as follows:
$\|V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}\|\leq\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$ (21)
Proof.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||$
$\leq||V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi^{*}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||+||V_{\mathcal{M}^{\prime}\cup R}^{\hat{\pi}^{*}}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||$
@@
Appendix D Implementation Details
We used two identical 2-layer MLP with 100 hidden units and ReLU activiations for the reward term $r$ and the shaping term $\phi$. The neural nets are trained using Adam optimizer (Kingma, 2014) with learning rate being $0.0003$. For the learned dynamic model, we also used a 2-layer MLP with ReLU activations and 128 hidden units. Gradient clipping is added on the model network with the value of 10. The model network is trained by the Adam optimizer as well with learning rate $0.0003$. For SAC and PPO, we follow the implementation from CleanRL repository (Huang et al., 2022).
Appendix E Graph Results
Below is the testing return diagrams from both deterministic and stochastic Mujoco Environments
Figure 2: Training return diagram of benchmarks and our algorithm in the InvertedPendulum-v4. We only show the training return until reaching the experts performance of all the other benchmarks, but the thorough training return of our algorithm.
Figure 3: Training return diagram of benchmarks and our algorithm in the InvertedDoublePendulum-v4.
Figure 4: Training return diagram of benchmarks and our algorithm in the Hopper-v3.

### v2

@@ Appendix D Performance Gap Analysis
Lemma D.2 (Explicit Feasible Reward Function (metelli2021provably)).
@@
With  Equation 7, we can derive the following error bound between $R\in\mathcal{R}^{E}_{\mathfrak{B}}$ and $\hat{R}^{E}\in\mathcal{R}_{\hat{\mathfrak{B}}}$.
Theorem D.3 (Reward Function Error Bound).
@@
Proof D.4.
From  D.2, we can derive the following representations of $R$ and $\hat{R}$ with the same set of $V$ and $\xi$:
@@
Next, we will propagate this bound to the value functions of optimal policy regarding different reward functions $R^{E}$ and $\hat{R}^{E}$. From the traditional Value iteration, we can write out the value function.
$V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\mathcal{T}(s^{\prime}|s,a)\left[R^{E}(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s^{\prime})\right].$ (9)
Lemma D.5 (Value Function Error under same policy and different rewards and MDP).
$V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||$: the performance difference of the same policy in different MDPs.
$V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}^{E}}^{\pi}(s)||\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|.$ (10)
Proof D.6.
[…]
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}||\mathcal{T}(s^{\prime}|s,a)\left[R^{E}(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R^{E}(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s^{\prime})\right]$
$+\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R^{E}(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[\hat{R}^{E}(s,a)+\gamma V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}^{E}}^{\pi}(s^{\prime})\right]||$
[…]
@@
Lemma D.7.
Let $\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}^{E}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}^{E}}^{\pi_{2}}(s)\|$ denote the performance difference between different policies $\pi_{1}$ and $\pi_{2}$ in the same learned MDP (viano2021robust; zhang2020learning). The following inequality holds:
$\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}^{E}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}^{E}}^{\pi_{2}}(s)\|\leq\frac{\gamma}{(1-\gamma)^{2}}\epsilon_{\mathcal{T}}R_{\max}.$
Theorem D.8 (Performance Difference Bound).
The performance difference between the optimal policies ($\pi^{*}$ and $\hat{\pi}^{*}$) in corresponding MDPs ($\mathcal{M}^{\prime}\cup R^{E}$ and $\hat{\mathcal{M}}^{\prime}\cup\hat{R}^{E}$) can be bounded as follows:
$\|V_{\mathcal{M}^{\prime}\cup R^{E}}^{\pi^{*}}-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}^{E}}^{\hat{\pi}^{*}}\|\leq\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$ (12)
Proof D.9.
[…]
@@
Appendix E Ablation Studies
E.1 Robustness to stochasticity
[…]
Figure 2: Training return diagram averaging across three seeds for different numbers of expert trajectories in InvertedPendulum-v4.
Std Expert Ours
0.5 $_{\pm 0.0} 1000.0±0.0100$$_{\pm 0.0}$
0.55 $802.4_{\pm 305.8}$ $1000.0_{\pm 0.0}$
0.6 $582.1_{\pm 360.5}$ $906.3_{\pm 59.1}$
0.65 $438.2_{\pm 322.3}$ $709.9_{\pm 80.6}$
0.7 $270.7_{\pm 236.0}$ $472.7_{\pm 85.7}$
Table 5: Best performance of expert and our method in InvertedPendulum-v4 environments with different Gaussian noises (standard deviations ranging from $0.5$ to $0.7$) for stochasticity under provided 100 expert trajectories.
E.2 Model estimation error and reward learning
In this study, we empirically evaluate the effect of dynamic model learning errors on our method’s performance, extending the theoretical analysis presented in subsection 4.3. To isolate the impact of model errors specifically on reward learning, we use SAC on real trajectories for policy optimization, thereby removing any influence of model errors on trajectory generation that would typically affect model-based policy optimization. To quantify the relationship between model errors and performance, we standardize the ensemble model architecture as 2-layer MLPs with varying hidden layer dimensions from $8$ to $256$ to adjust model capacity. Our experiments are conducted in HalfCheetah-v4 with random, policy-unknown Gaussian noise (mean 0 and standard deviation $0.5$), as described in section 5. From Figure 4 and Figure 3, we observe the general trend that as modeling error decrease together with increasing capacity of the model structure, performances also increases, which is obvious when hidden dimension bumps up from $8$ to $16$ and $16$ to $32$. As transition model error narrows down, the performance improvement also becomes less obvious.
Figure 3: Transition model learning error diagram averaging across three seeds for 10 expert trajectories in HalfCheetah-v4.
Figure 4: Training return diagram averaging across three seeds for 10 expert trajectories in HalfCheetah-v4.
E.3 Does model-based trajectories generation help?
[…]
Figure 5: Performance diagram averaging across three seeds for different algorithms in Hopper-v4 with 1000 expert trajectories provided. DAC is in green color; mbirl_sacmbirl\_sac is in blue; Our method is in red.
Method Performance
DAC $2007.1_{\pm 719.7}$
mbirl_sacmbirl\_sac $2694.5_{\pm 77.5}$
Ours $2798.8_{\pm 82.9}$
Table 6: Best performance of three methods in stochastic Hopper-v4 environment with under provided 1000 expert trajectories.
Appendix F Algorithmic Framework
Algorithm 1 Model-Enhanced IRL (ME-IRL)
1:Input: Expert buffer $\mathcal{D}_{exp}$ (collected expert trajectories), total training steps NN, starting step for training policy STARTING_STEP, horizon length $H$ for synthetic trajectory generation.
2:Output: Learned policy $\pi$ optimized to mimic expert behavior.
3: Initialize policy $\pi$ (random or pre-trained), discriminator $D_{\theta}$, buffers $\mathcal{D}_{env}$ (real environment samples), $\mathcal{D}_{gen}$ (synthetic samples), and transition model $\hat{\mathcal{T}}$ (random initialization).
4:for step $t$ in {1, …, N} do
5:  Interact with the environment using policy $\pi$ to collect new state-action pairs.
6:  Add the collected state-action pairs $(s_{t},a_{t},s_{t+1})$ to the real environment buffer $\mathcal{D}_{env}$.
7:  if $t<\text{STARTING\_STEP}$ then
8:   Pre-train transition model: Train $\hat{\mathcal{T}}$ on batches of samples from $\mathcal{D}_{env}$ using maximum likelihood estimation (MLE) loss to stabilize the transition model early.
9:  else
10:   Step 1: Update Transition Model and Generate Synthetic Data
11:   Update the transition model $\hat{\mathcal{T}}$ using maximum likelihood loss (MLE) on $\mathcal{D}_{env}$.
12:   Use $\hat{\mathcal{T}}$ to generate $H$-step synthetic trajectories and store them in the synthetic buffer $\mathcal{D}_{gen}$.
13:   Step 2: Update Discriminator
14:   Sample mini-batches of state-action pairs from $\mathcal{D}_{exp}$ (expert buffer) and $\mathcal{D}_{env}$ (environment buffer) with varying ratio.
15:   Train the discriminator $D_{\theta}$ using cross-entropy loss to classify expert data (from $\mathcal{D}_{exp}$) versus policy-generated data (from $\mathcal{D}_{env}$) as described in  Equation 3.
16:   Step 3: Policy Optimization with Mixed Data
17:   Sample state-action batches from $\mathcal{D}_{env}$ and $\mathcal{D}_{gen}$ with a varying ratio that increases the use of $\mathcal{D}_{gen}$ as $\hat{\mathcal{T}}$ becomes more accurate.
18:   Compute the reward $\hat{R}_{\theta}$ for the sampled state-action pairs using the discriminator.
19:   Update the policy $\pi$ using Soft Actor-Critic (SAC) with the computed reward $\hat{R}_{\theta}$ as the optimization objective.
20:  end if
21:end for
22:Return: Optimized policy $\pi$.
Appendix G Implementation Details
For our framework, we use two identical 2-layer Multi-Layer Perceptrons (MLPs) with 100 hidden units and ReLU activations for both the reward function $R$ and the shaping potential function $\phi$. To initialize the replay buffer for both DAC and ours, we collect 1,000 steps samples in InvertedPendulum-v4 and InvertedDoublePendulum-v4, and 10,000 steps samples in Hopper-v4, HalfCheetah-v4, and Walker2d-v4 with initial policy. During this pre-training phase, we also update the transition model at each step to mitigate divergence might happen at the beginning of the training. Additionally, the transition model is only trained using samples from real environment buffer $\mathcal{D}_{env}$ in policy optimization section before actor and critics updates during training phase. As discussed in section 4, the size of the synthetic data buffer $\mathcal{D}_{\text{gen}}$ and the ratio of samples drawn from it increase as the model accuracy improves. Both parameters increase linearly with training steps, up to a maximum synthetic-to-real data ratio of 0.5 per training step and a maximum buffer size of $1$ million samples in $\mathcal{D}_{\text{gen}}$. For consistency in comparisons, we used similar network structures and hyper-parameters for AIRL, GAIL, and DAC baselines, which we reference the implementations from arulkumaran2024pragmatic and gleave2022imitation. Detailed hyper-parameters for these networks are provided in the table below. For on-policy baselines AIRL and GAIL, the rollout length is set to 1,000 for InvertedPendulum-v4 and InvertedDoublePendulum-v4, and 5,000 for Hopper-v4, Walker2d-v4, and HalfCheetah-v4. For the SAC and PPO policy optimization components, we reference implementations from the CleanRL repository (huang2022cleanrl). Our implementation of Dreamer-v2 and RSSM are based on SheepRL (EclecticSheep_and_Angioni_SheepRL_2023).
Table 7: Hyper-parameters table.
Hyper-parameter Value
Seeds 0, 5, 10
Buffer Size 1M
Batch Size 128
Max Grad Norm 10
Starting Steps 1,000/10,000
Global Timesteps 100k/1M
Discount Factor 0.99
Model-based Policy Optimization
Learning Rate for Actor 3e-4
Learning Rate for Critic 3e-4
Learning Rate for Model 3e-4
Network Layers 3
Policy Network Neurons [64, 64]
Critic Network Neurons [128, 128]
Model Network Neurons [256, 256]
Ensemble Numbers 7
Elites Numbers 5
Activation Tanh(Policy)/ReLU
Optimizer Adam
Initial Entropy $|\mathcal{A}|$
Learning Rate for Entropy 3e-4
Train Frequency for Actor 1
Train Frequency for Critic 1
Train Frequency for Model 1
Synthetic and Real Data Mix Coef 0.5
Horizon($H$) 4
Adversarial Discriminator
Learning Rate 3e-4
$R$ Network Neurons [100, 100]
$\phi$ Network Neurons [100, 100]
Optimizer Adam
Loss Binary Cross-Entropy
Appendix H MuJuCo Graphical Results
Below are the testing return diagrams from stochastic MuJoCo Environments. Since AIRL and GAIL use distinct environmental training steps from DAC and our method,  Figure 6 provides a clear comparison under 10M landscape while the rest of the graphs show the sample efficiencies for all algorithms under 1M landscape.
Figure 6: Training curves of all 4 methods in 5 different stochastic environments with 100 expert trajectories. For better comparison in sample efficiency, graph is presented under 10M landscape.
Figure 7: Training return diagram averaging across three seeds for different numbers of expert trajectories in Stochastic InvertedPendulum-v4.
Figure 8: Training return diagram averaging across three seeds for different numbers of expert trajectories in Stochastic InvertedDoublePendulum-v4.
Figure 9: Training return diagram averaging across three seeds for different numbers of expert trajectories in Stochastic Hopper-v4.
Figure 10: Training return diagram averaging across three seeds for different numbers of expert trajectories in Stochastic HalfCheetah-v4.
Figure 11: Training return diagram averaging across three seeds for different numbers of expert trajectories in Stochsatic Walker2d-v4.
Appendix I Supplementary Material
I.1 Codespace
To help make our results reproducible, we provide our implementation at the following anonymous link: https://anonymous.4open.science/r/MBIRL-4C2F/README.md
Method Retern
Expert $22300.0_{\pm 3662.0}$
CNN-AIRL $1200.0_{\pm 748.3}$
Ours $14400.0_{\pm 5607.1}$
Table 8: Mean return over 10 evaluation episodes after 400k steps on BattleZone-v5.
I.2 Atari Experiment Details
In this section, we explain the implementation details for the Atari experiments. We follow SheepRL (EclecticSheep_and_Angioni_SheepRL_2023) for our implementation for the expert policy (Dreamer-v2 (hafner2020mastering)) and the baseline IRL comparison (CNN-AIRL(tucker2018inverse)). As for our own algorithm, we replace the ensemble dynamic model with RSSMhafner2019dream to capture the high-dimensional inputs from Atari environments. To address the discrete action space, we use categorical distributions as actor networks for all algorithms and change the policy update logic accordingly. Due to computation resource constraint, we test our algorithms on two distinct environments - SpaceInvaders-v5 and BattleZone-v5 instead of the full set of Atari2600 environments. All algorithms are trained for 400k steps. We report the mean return for SpaceInvaders-v5 in Table 4 and BattleZone-v5 in Table 8. The hyper-parameters of the RSSM can be found in Table 9.
Table 9: RSSM hyper-parameters for Atari experiments.
Policy Optimization for discrete actor
Discrete Latent Size 32
Stochastic Latent Size 32
KL Balancing $\alpha$ 0.8
Free Nats 1.0 (averaged)
KL Regulariser Scale 1.0
Encoder / Observation Model
CNN Channel Multiplier 48
MLP Layers 4
Dense Units 400
Activation ELU / ELU
Layer Norm False
Recurrent Model
Recurrent State Size 600
Dense Units 400
Activation ELU
Layer Norm True
Transition & Representation Models
Hidden Size 600
Activation ELU
Layer Norm False
Discount Model
Dense Units 400
Activation ELU / ELU
Layer Norm False